OpenStudy (anonymous):

Find the y value for the point that splits segment EF in half if point E is located at (5, −2) and point F is located at (−1, 3). 1 0.5 4.5 2

OpenStudy (imstuck):

You are looking for the midpoint. Do you know the formula for that?

OpenStudy (imstuck):

\[midpt = (\frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{2 })\]

OpenStudy (anonymous):

I don't know sorry...:/ I'm clueless

OpenStudy (imstuck):

ok, use your points and fill it in like this:

OpenStudy (imstuck):

\[(\frac{ 5+(-1) }{2 },\frac{ -2+3 }{ 2})\]

OpenStudy (anonymous):

would it be 4.5?

OpenStudy (imstuck):

-2 + 3 is 1, and you still have to divide it by 2. So it would be .5.

OpenStudy (anonymous):

I think I get it now can we try another problem together?

OpenStudy (imstuck):

The midpoint is (2, 1/2).

OpenStudy (imstuck):

Yes, we can do another.

OpenStudy (anonymous):

Find the perimeter of the following shape, rounded to the nearest tenth: 12.5 15.7 16 16.5

OpenStudy (anonymous):

Thank you by the way

OpenStudy (imstuck):

You're welcome btw!

OpenStudy (imstuck):

You need to find the length of one short side and one long side. Then double them, cuz you have a rectangle, and a rectangle has two short sides the same length and two long sides the same length. so we got lucky here and need to only find the length on one short and one long.

OpenStudy (imstuck):

The distance formula needs to be used here. Do you know it?

OpenStudy (anonymous):

I don't know it :/

OpenStudy (imstuck):

That's ok. You're learning a lot today!

OpenStudy (anonymous):

Okay and that's good I need it

OpenStudy (imstuck):

\[d=\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\]

OpenStudy (imstuck):

Your points for a short side are A and B and the coordinates for A are (5,5) and B (7,3).

OpenStudy (anonymous):

for perimeter u add right? Or times

OpenStudy (imstuck):

The distance formula looks like this with your points:\[d=\sqrt{(7-5)^{2}+(3-5)^{2}}\]which translates to \[d=\sqrt{(2)^{2}+(-2)^{2}}\]

OpenStudy (anonymous):

okay that's easy

OpenStudy (imstuck):

perimeter is addding, btw.

OpenStudy (imstuck):

The length of AB is \[AB=\sqrt{8}\]

OpenStudy (imstuck):

Now let's do BC in the same way.

OpenStudy (anonymous):

0,3 and 3,7

OpenStudy (imstuck):

\[d=\sqrt{(3-7)^{2}+(0-3)^{2}}\]or\[d=\sqrt{(-4)^{2}+(-3)^{2}}\]or \[d=\sqrt{25}\]and d = 5 for the long side.

OpenStudy (imstuck):

So here's your rectangle:

OpenStudy (imstuck):

Perimeter is 5 + 5 + sqrt 8 + sqrt 8

OpenStudy (anonymous):

16.10

OpenStudy (imstuck):

Are you supposed to find the exact value or are you supposed to leave a radical sign in the answer?

OpenStudy (imstuck):

Oh I see by the answer choices that you multiply it in. Ok, got that much. Hold on.

OpenStudy (imstuck):

Do you see the answer there? It is there, rounded to the nearest tenth.

OpenStudy (imstuck):

It's not 16.1 though.

OpenStudy (anonymous):

is it 16

OpenStudy (imstuck):

\[\sqrt{8}=2.828427125\]You have two sides that length so multiply that by 2 to get 5.656854249

OpenStudy (imstuck):

Now add the 2 sides that measure 5 each (10 total) to that 5.656854249 and get 15.65685425, which is 15.7 rounded to the nearest tenth.

OpenStudy (anonymous):

I got 16 when I added it then I rounded it to the nearest tenth

OpenStudy (imstuck):

You couldn't have gotten an even 16 because the square root of 8 is an irrational number and the decimal goes on forever. You would HAVE to have a decimal if you did it right.

OpenStudy (imstuck):

Look back over the work I showed for you to see how I did it. I typed it all out step-by-step. I have to go to work now. ok?

OpenStudy (imstuck):

The answer is 15.7...promise.

OpenStudy (anonymous):

oh I think I added it wrong and okay thank you!

OpenStudy (imstuck):

You are so welcome!

OpenStudy (anonymous):

when u get back I have 3 more questions I need help with..

OpenStudy (anonymous):

The distance formula can be used to prove a triangle has congruent sides a right angle parallel sides congruent angles

OpenStudy (anonymous):

When the coordinates (2, 3), (4, 4), (6, 3), and (4, 2) are joined, which shape is formed? Parallelogram Rectangle Rhombus Square

OpenStudy (anonymous):

Line AB contains points A (0, 1) and B (1, 5). The slope of line AB is −4 negative 1 over 4 1 over 4 4

OpenStudy (anonymous):

@IMStuck please help me again